# Exotic Superfluid with Emergent Flux in a One-Dimensional Bose-Fermi Mixture

### v3

@@ Abstract
We find a novel chiral superfluid (CSF) phase in a chain of Bose-Fermi mixture, which has been validated using two unbiased numerical methods, density matrix renormalization group and Grassmann multi-scale entanglement renormalization ansatz. The system hosts the interplay between two types of fermions: bare spinless fermions and composite fermions, the latter consisting of a fermion and a boson. In the CSF phase, bosons condensate at non-zero momentum ±$\pm 2\pi/L$ with chain length $L$. The local superfluid order parameter continuously rotates along the chain, indicating that the CSF phase spontaneously breaks time-reversal symmetry. This symmetry breaking gives rise to an emergent flux in the background, effectively optimizing the kinetic energy of the composite fermions within the system. We provide a physical understanding at the mean-field level. Furthermore, we demonstrate that the 1D CSF phase can emerge in a more widely applicable extended Hubbard model. The potential realization of this phase in cold-atom experiments has also been explored.
@@
The interaction between bosons and fermions is essential in condensed matter physics and can give rise to exotic states of matter. In conventional superconductors, phonon exchange between electrons triggers the formation of Cooper pairsSchrieffer (1983). Over the last two decades, ultracold bosonic and fermionic atoms have emerged as remarkable platforms for simulating quantum physicsBloch et al. (2008); Gross and Bloch (2017); Schfer et al. (2020). Due to their high controllability, atomic mixtures have exhibited rich phenomena, such as double superfluidsFerrier-Barbut et al. (2014); Roy et al. (2017); Yao et al. (2016), mediated interactionsDeSalvo et al. (2019), heteronuclear moleculesInouye et al. (2004); Stan et al. (2004); Milczewski and Duda (2023), and hybridized sound modesPatel et al. (2023); Shen et al. (2024). Furthermore, topological $p$-wave superfluids in BFMs have garnered theoretical attentionWu and Bruun (2016); Kinnunen et al. (2018b). Remarkably, spontaneous translational symmetry-breaking insulating states have recently been observed in two distinct solid-state BFM systemsRuiz-Tijerina (2023); Zeng et al. (2023); Lagoin et al. (2023). Consequently, BFMs are fertile ground for cultivating intriguing new superfluid phases.
In this paper, we reveal a novel chiral superfluid (CSF) phase that spontaneously breaks time-reversal symmetry. In this phase, the bosonic superfluid order rotates along the chain, resulting in an effective emergent flux. The system we investigate exhibits a complex interplay between bare spinless fermions and composite fermions (consisting of a fermion and a boson) in a 1D BFM. In the presence of composite hopping, where a fermion hops with a boson, the tunneling processes of fermions and bosons become correlated. Remarkably, the system solely adopts the CSF as its ground state when an even number of fermions are present in the mixture. Conversely, it transitions to a conventional superfluid state when the number of fermions is odd. This indicates that the chirality of this quantum system, protected by an energy gap, can be manipulated by adding or removing particles through the tuning of chemical potentials. This even-odd disparity underscores the topological nature of the CSF. The CSF phase can be directly realized within the conventional extended Bose-Fermi Hubbard model. We provide numerical evidence and discuss the possibility of realizing the 1D CSF in cold-atom experiments.
Emergence of CSF phase.— We introduce a minimal model that incorporates the motion of composite fermions (CFs) $-t_{cf}c^{\dagger}_{i}b^{\dagger}_{i}c_{j}b_{j}$, which describes the binding and hopping of a fermion and a boson. The resulting Hamiltonian for our system reads
$H=$ $-t_{f}\sum_{\langle ij\rangle}(c^{\dagger}_{i}c_{j}+H.c.)-t_{cf}\sum_{\langle ij\rangle}(c^{\dagger}_{i}b^{\dagger}_{i}c_{j}b_{j}+H.c.)$ (1)
$+\frac{U_{bb}}{2}\sum_{i}n^{b}_{i}(n^{b}_{i}-1)-\mu_{b}\sum_{i}n^{b}_{i}-\mu_{f}\sum_{i}n^{f}_{i}$
where $U_{bb}$ represents the onsite repulsive interaction between bosons, $n^{b}_{i}(n^{f}_{i})$ is the onsite boson (fermion) number operator, and $\mu_{b}(\mu_{f})$ is the chemical potential of bosons (fermions). We consider the case $N_{f}>N_{b}$.
[…]
We address the origin of CSF from two distinct perspectives, starting from the bare fermion limit ($t_{f}\gg t_{cf}$). First, the kinetic term of composite particles ($t_{cf}$) suggests that a bosonic superfluid in the background induces projected ”silhouette” composite particles. A CSF introduces an additional phase factor that fermions acquire during hopping. Effectively, this generates an ”emergent flux” that further improves the system energy, analogous to the behavior of electrons hopping in a ring with a magnetic fluxPeter et al. (1989). Based on this argument, we develop a mean-field consideration, resulting in an effective boson-free model that qualitatively aligns with our original Hamiltonian in the bare fermion limit. Second, the fermionic response to the bosonic emergent flux is crucial. As the CSF emerges, it profoundly modifies the fermionic momentum distribution. Numerical calculations reveal that the degeneracy of the fermionic Fermi surface (two Fermi points) is lifted simultaneously when the time-reversal symmetry of the superfluid is broken. This lifting of degeneracy is accompanied by a redistribution of the momentum distribution of composite particles, which becomes more concentrated around small $k$ values, further optimizing the system’s energy. The collaborative effect between the fermionic and bosonic sectors stabilizes the CSF state.
@@
Figure 1: (a)–(c):The boson superfluid order parameter $\langle b_{i}\rangle$ (arrows) measured on each site $i$ (open circles), magnified ten times, with the horizontal/vertical component of the arrow corresponding to its real/imaginary part, respectively. (d): The argument angle $\theta_{i}$ of the nearest-neighbor hopping term $\langle b_{i}^{\dagger}b_{i+1}\rangle$. Results are obtained at parameter setting $t_{f}=0.2,t_{cf}=0.4,U_{bb}=3.0.$
@@
CSF and Fermions parity.— Furthermore, our system exhibits an intriguing sensitivity to the parity of fermions. Specifically, when the number of fermions $N_{f}$ is even, the ground state is in the CSF phase where bosons condensation occurs at $k_{0}=\pm 2\pi/L$, either clockwise or counterclockwise, depending on the sign of $k_{0}$. However, when $N_{f}$ is odd, bosons condense at $k=0$, similar to a conventional superfluid. In both cases, the superfluid phase persists to an intense competition region $t_{f}\approx t_{cf}$. Numerical results from the long-range boson correlation $G_{b}(x)=\langle b_{i}^{\dagger}b_{i+x}\rangle$ and momentum distribution $n_{b}(k)$ show that superfluidity emerges as soon as $t_{cf}$ is turned on, such as ($t_{cf}=10^{-3}t_{f}$). In the weakly coupled case, with a boson superfluid as the background, we understand that the free hopping fermions can gain additional energy through the $t_{cf}$ term in Eqn. 1, by projecting out silhouette composite particles. When $N_{f}$ is even, the energy contribution from the $t_{cf}$ term can be further optimized by allowing superfluid order parameter $\langle b\rangle$ to rotate and acquire chirality. In such cases, the hopping amplitude $\langle b_{i}^{\dagger}b_{j}\rangle$ becomes a complex number due to the phase difference of $\langle b\rangle$ between neighboring sites. Consequently, the expression $c_{i}^{\dagger}c_{j}b_{i}^{\dagger}b_{j}$ should be regarded as $c_{i}^{\dagger}c_{j}\langle b_{i}^{\dagger}b_{j}\rangle$. Numerical results confirm that the measurement of $\langle b_{i}^{\dagger}b_{j}\rangle$ on different bond $\langle i,j\rangle$ exhibit nearly identical arguments and moduli, as shown in Fig. 1(d). This observation suggests that a CSF generates an additional phase factor $e^{i\theta_{ij}}$ in the composite hopping term.
@@
Figure 2: The ground state energy $E(\theta)$ of the mean-field Hamiltonian $H_{MF}$ as a function of the argument $\theta$ in $L=18,36,54$ systems when $t_{f}=0.2,t^{\prime}_{cf}=0.4$. (a) Even fermion parity. The fermion filling factor is fixed at $n_{f}=N_{f}/L=4/9$. There is a prominent peak on $\theta=0$ where no magnetic flux is threaded through the system. And $E(\theta)$ reaches its absolute minimum at $\theta_{opt}=\pi/L$. (b) Odd fermion parity. $N_{f}$ are $7,15,23$ for $L=18,36,54$ respectively. No peak shows on $\theta=0$ contrary to the even case in the main panel. And $E(\theta)$ reaches its absolute minimum at $\theta_{opt}=0$
@@
Figure 3: Momentum distribution functions $n(k)$ for $L=36$ system with boson filling factor $n_{b}\approx 1/4$, obtained by GMERA. ”FS” refers to the symmetric Fermi surface of original fermions. (a) For $N_{f}=16$ (even), bosons condensate at $|k_{0}|=2\pi/L$ (position of the blue peak in $n_{b}(k)$). $k^{c}_{f}=\sum_{k}n_{f}(k)\times k/N_{f}=-0.0122\pi$ is the center of original Fermi sea, whereas $k^{c}_{cf}=\sum_{k}n_{cf}(k)\times k/N_{cf}=0.0075\pi$ is the center of Fermi sea for composite fermion, where $N_{cf}=\sum_{k}n_{cf}(k)$. (b) For $N_{f}=15$ (odd), bosons condensate at $k=0$. The parameters are chosen as $t_{f}=0.2,t_{cf}=0.4,U_{bb}=3.0.$ Fourier transforms of $G_{b}(i,j)=\langle b_{i}^{\dagger}b_{j}\rangle$, $G_{f}(i,j)=\langle c_{i}^{\dagger}c_{j}\rangle$, $G_{cf}(i,j)=\langle c_{i}^{\dagger}b_{i}^{\dagger}c_{j}b_{j}\rangle$ and $G_{hf}(i,j)=\langle c_{i}^{\dagger}b_{i}c_{j}b_{j}^{\dagger}\rangle$ give respective momentum distributions $n_{b}(k),n_{f}(k),n_{cf}(k)$ and $n_{hf}(k)$. The black dashed lines guide the eyes at $k=0$.
@@
$H_{MF}=-t_{f}\sum_{\left\langle ij\right\rangle}(c^{\dagger}_{i}c_{j}+H.c.)-t^{\prime}_{cf}\sum_{\left\langle ij\right\rangle}(c^{\dagger}_{i}c_{j}e^{i\theta_{ij}}+H.c.)$
@@
The second term can be intuitively understood through our mean-field ansatz, treating the influence of bosons classically as a pure phase factor $e^{i\theta_{ij}}$. Therefore, we substitute the hopping term $c_{i}^{\dagger}b_{i}c_{i+1}b_{i+1}^{\dagger}$ with $c_{i}^{\dagger}c_{j}e^{-i\theta}$. This model highlights the two competing hopping processes: the former stands for original free hopping, whereas the latter depicts hoppings experiencing a lattice flux. This simple model can be solved straightforwardly. As shown in Fig. 2, the ground state energy $E$ exhibits a periodically varying function superimposed on a parabolic background, illustrating the competition between two distinct hopping processes. Remarkably, a significant difference is also observed between systems with even (left panel) and odd (right panel) fermion parity: when $N_{f}$ is even, the optimal $\theta$ equals to $\pi/L$. In contrast, when $N_{f}$ is odd, the optimal $\theta$ takes the value of 00. This observation from the mean-field model is similar to the results obtained in the BFM. However, it is crucial to note that $\theta_{ij}$, the argument of the complex $\langle b_{i}^{\dagger}b_{i+1}\rangle$ shown in Fig. 1(d), averages out to $0.0940\pi$ for $L=18$ or $-0.0491\pi$ for $L=36$, deviating from the optimal angle of ±$\pm\pi/L$ predicted by mean-field theory. This deviation is attributed to the periodic boundary conditions imposed on our homogeneous lattice.
@@
Shift of Fermi surface.— Interplay and correlations between different species of particles lead to the shift of the Fermi surface. We focus on the correlation function and momentum distribution of involved particles, influenced by the effective flux. When $N_{f}$ is odd, the conventional bosonic superfluid serves as a flux-free background, the correlation function of projected composite fermion $G_{cf}(i,j)=\langle c_{i}^{\dagger}b_{i}^{\dagger}c_{j}b_{j}\rangle$ shows similar long-range behavior as that of an original fermion. As demonstrated in Fig. 3, two clear Fermi surfaces (points) of composite particles have the same momentum as those of spinless fermions. In contrast, when $N_{f}$ is even, the ground state exhibits a two-fold degeneracy: bosons condensate at either $+2\pi/L$ or $-2\pi/L$, time-reversal symmetry of the system is spontaneously broken. Intriguingly, this degeneracy lift simultaneously affects the Fermi surface of a bare fermion. The selection of the last filled momentum is determined by the direction of the effective flux, resulting in an asymmetric bulk of the Fermi surface. Consequently, the Fermi surface of the projected composite fermion shifts towards $k=0$, optimizing the system’s overall energy. In this uncorrelated limit, this projection can be rationalized as follows: $n_{cf}(k)=1/L\sum_{i,j}G_{cf}(i,j)e^{ik(x_{i}-x_{j})}\approx 1/L\sum_{i,j}G_{f}(i,j)G_{b}(i,j)e^{ik(x_{i}-x_{j})}\approx 1/L\sum_{i,j}G_{f}(i,j)|A|e^{i(k+k_{0})(x_{i}-x_{j})}.$ Here, $G_{b}(i,j)=|A|e^{ik_{0}(x_{i}-x_{j})}$ represents the boson correlation for superfluid condensates at momentum $k_{0}$.
@@
Figure 4: The entanglement entropy $S_{vN}$ as a function of $L$ at different values of $t_{cf}/t_{f}$ and fermion parity. Boson and fermion numbers are $N_{f}=N_{b}=L/2$ in ’a’ (blue solid squares) and ’b’ (red solid circles), $N_{f}=L/2-1,N_{b}=L/2$ in ’c’ (orange solid diamonds) and $N_{f}=N_{b}=L/2$ in ’d’ (green solid triangles). Inset shows the results of central charge.
@@
Figure 5: Momentum distribution functions $n(k)$ obtained by DMRG. (a) $N_{f}=16=2N_{b}$ and (b)$N_{f}=15,N_{b}=8$ correspond to the Hamiltonian (1) at $t_{cf}=2t_{f}=0.4$ for $L=36$ sites. (c)$N_{f}=6,N_{b}=3$ and (b)$N_{f}=7,N_{b}=4$ correspond to the extended BF Hamiltonian (Exotic Superfluid with Emergent Flux in a One-Dimensional Bose-Fermi Mixture) at $t_{f}=0.4,t_{b}=0.05,U_{bb}=3.0,U_{bf}=3.0,V_{bf}=-0.3$ for $L=18$ sites. For even $N_{f}$ (Left column), bosons condensate at $|k_{0}|=2\pi/L$ (positions of the two blue peaks in $n_{b}(k)$). For odd $N_{f}$ (right column), bosons condensate at $k=0$.

### v4

@@ Abstract
We find a novel chiral superfluid (CSF) phase in a one-dimensional Bose-Fermi Hubbard model with significant mass and density imbalance between the two species. In the CSF phase, bosons condensate at non-zero momentum ±$\pm 2\pi/L$ with chain length $L$. To capture the essential physics of this new phenomenon, we study an alternative simplified model that only features competition between single-fermion hopping and hopping of composite particles composed of a fermion and a boson. This model captures the low energy physics of the Hubbard model and hosts a robust CSF phase. Our unbiased numerical studies show that in the CSF phase, the local superfluid order parameter continuously rotates along the chain, indicating that time-reversal symmetry is spontaneously broken. This symmetry breaking generates an emergent flux in the background, effectively optimizing the system’s ground-state energy. We provide a physical understanding at the mean-field level. Furthermore, we have explored the potential realization of this phase in cold-atom experiments.
@@
The interaction between bosons and fermions is essential in condensed matter physics and can give rise to exotic states of matter. In conventional superconductors, phonon exchange between electrons triggers the formation of Cooper pairsSchrieffer (1983). Over the last two decades, ultracold bosonic and fermionic atoms have emerged as remarkable platforms for simulating quantum many-body physicsBloch et al. (2008); Gross and Bloch (2017); Schfer et al. (2020). Due to their high controllability, atomic mixtures have exhibited rich phenomena, such as double superfluidsFerrier-Barbut et al. (2014); Roy et al. (2017); Yao et al. (2016), mediated interactionsDeSalvo et al. (2019), heteronuclear moleculesStan et al. (2004); Milczewski and Duda (2023); Duda et al. (2023), and hybridized sound modesPatel et al. (2023); Shen et al. (2024). Furthermore, topological $p$-wave superfluids in BFMs have garnered theoretical attentionWu and Bruun (2016); Kinnunen et al. (2018b). Remarkably, spontaneous translational symmetry-breaking insulating states have recently been observed in two distinct solid-state BFM systemsRuiz-Tijerina (2023); Zeng et al. (2023); Lagoin et al. (2023). Consequently, BFMs are fertile ground for cultivating intriguing new superfluid phases.
In this paper, we reveal a novel chiral superfluid (CSF) phase that spontaneously breaks time-reversal symmetry. In this phase, the bosonic superfluid order rotates along the one-dimensional (1D) periodic chain, creating an effective emergent flux. Consequently, bosons condense at non-zero momentum ±$\pm 2\pi/L$, where $L$ denotes the chain length. Initially, we find the CSF within the extensively studied Bose-Fermi Hubbard (BFH) model Sowiński and García-March (2019); Pollet et al. (2006); Lewenstein et al. (2004), which describes a mixture of ultracold bosons and fermions confined within an optical lattice. Our investigation of the BFH model focuses on conditions characterized by a substantial imbalance in both the masses and densities of the two species, a regime that has received limited attention in previous studiesSowiński and García-March (2019); Pollet et al. (2006); Lewenstein et al. (2004); Rizzi and Imambekov (2008); Fratini and Pieri (2012); Mering and Fleischhauer (2008). Due to strong boson-fermion interactions, the low-energy physics of the system can be described by an effective model where composite-particle hopping competes with single-fermion hopping. Composite-particle hopping was introduced in the context of high-temperature superconductivitySheshadri and Chainani (2020). In our study, composite-particle hopping (a fermion hops together with a boson) captures the essence of boson-fermion correlations, which directly leads to the emergence of CSF. Based on this effective model, we provide a comprehensive understanding of the origin of CSF from two distinct perspectives, supported by concrete numerical evidence and mean-field analysis.
Remarkably, our system solely adopts the CSF as its ground state when an even number of fermions are present in the mixture. In contrast, it behaves as a normal superfluid state when the number of fermions is odd. This indicates that the chirality of this quantum system, protected by an energy gap, can be manipulated by adding or removing particles through the tuning of chemical potentials. This even-odd disparity underscores the topological nature of the CSF. Finally, we discuss the potential for realizing and detecting the 1D CSF in cold-atom experiments.
Figure 1: Momentum distribution functions $n(k)$ obtained by DMRG for the Hubbard model (upper panels) and the effective model (lower panels) shows similar behavior. (a) $N_{f}=6=2N_{b}$ and (b)$N_{f}=5,N_{b}=3$ correspond to the BFH Hamiltonian(1) at $t_{b}=0.25t_{f},U_{bf}=-2.5t_{f}$. (c)$N_{f}=6=2N_{b}$ and (d)$N_{f}=5,N_{b}=3$ correspond to the composite-particle Hamiltonian(2) at $t_{cf}=t_{f}$. For even $N_{f}$ (Left column), bosons condensate at $|k|=2\pi/L$ (positions of the two red peaks in $n_{b}(k)$). For odd $N_{f}$ (right column), bosons condensate at $k=0$. We set $U_{bb}=7.5t_{f}$ and $L=36$ for all cases.
CSF phase in the BFH Model.— The standard BFH model, featuring an onsite boson-fermion interaction strength $U_{bf}$ within a 1D periodic chain, is expressed as follows:
$H_{BFH}=$ $-t_{f}\sum_{i}(c^{\dagger}_{i}c_{i+1}+H.c.)-t_{b}\sum_{i}(b^{\dagger}_{i}b_{i+1}+H.c.)$ (1)
$+\frac{U_{bb}}{2}\sum_{i}n^{b}_{i}(n^{b}_{i}-1)+U_{bf}\sum_{i}n^{b}_{i}n^{f}_{i}$
where $c^{\dagger}_{i}$ ($b^{\dagger}_{i}$) creates a spinless fermion (boson) at site $i$, $t_{f}$ ($t_{b}$) is the tunneling amplitudes of fermions (bosons), and $n^{b}_{i}=b^{\dagger}_{i}b_{i}$ $(n^{f}_{i}=c^{\dagger}_{i}c_{i})$ denotes the boson (fermion) number operator. The term $U_{bb}$ suppresses double occupation of bosons and we set $U_{bb}=7.5t_{f}$ if not specified. We employ both the density matrix renormalization group (DMRG)White (1992); Fishman et al. (2022) and Grassmann multi-scale entanglement renormalization ansatz (GMERA)Lou and Chen (2015) methods, two well-established unbiased numerical techniques based on tensor product ansatz to conduct calculations. Our primary focus is on the attractive interspecies coupling ($U_{bf}<0$), though similar results are obtained on the repulsive side SM .
When the interspecies interaction is sufficiently strong, a composite fermion (CF) is formed through the pairing of a fermion and a boson, assuming equal densities of both species Pollet et al. (2006); Lewenstein et al. (2004); Rizzi and Imambekov (2008); Fratini and Pieri (2012); Powell et al. (2005). To introduce a density imbalance, we set the total number of fermions $N_{f}$ to be greater than the number of bosons $N_{b}$. This imbalance results in an excess of fermions without corresponding bosons to pair with, leading to strong scattering between fermions and composite particles and enhancing bosonic superfluidity.
To avoid the scenario where bosons do not exhibit condensation, we ensure that the fermion density remains below 1/3 Rizzi and Imambekov (2008). Additionally, we incorporate significantly unequal hopping amplitudes, reflecting the inherent mass difference between heavy bosons and light fermions. In the case of weak attraction, bosons occupy the normal superfluid phase with zero condensation momentum, while fermions are in the Luttinger liquid phase with a well-defined Fermi momentum. However, as the attractive strength increases, DMRG results indicate that bosons condense into the CSF phase at momentum $k=\pm 2\pi/L$ with even fermion number $N_{f}$ at $t_{b}=0.25t_{f}$, as illustrated in Fig. 1(a). In contrast, for odd $N_{f}$ in Fig. 1(b), bosons condensate at $k=0$, exhibiting conventional SF behavior. The distinct behavior between systems with even and odd parity recalls the behavior of fermions in a 1D periodic chain under the influence of magnetic flux Nakano (2000). This observation inspired us to propose a straightforward effective model that captures the essential physics of this mixed system.
@@
Figure 2: (a)–(c):The boson superfluid order parameter $\langle b_{i}\rangle$ (arrows) measured on each site $i$ (open circles), magnified ten times, with the horizontal/vertical component of the arrow corresponding to its real/imaginary part, respectively. (d): The argument angle $\theta_{i}$ of the nearest-neighbor hopping term $\langle b_{i}^{\dagger}b_{i+1}\rangle$, corresponding to cases (a)–(c). Results are obtained by GMERA on the composite-particle Hamiltonian(2) with parameters $t_{cf}=2t_{f},U_{bb}=7.5t_{f}$, and the following particle numbers: (a)$N_{f}=8=2N_{b}$ for $L=18$; ($N_{f}=16=2N_{b}$ for $L=36$; (c)$N_{f}=15,N_{b}=8$ for $L=36$.
Under the influence of boson-fermion attraction, a system characterized by an excess of fermions ($N_{f}>N_{b}$) can generally be viewed as a mixture of CFs and unpaired fermions. Whether bosons exhibit condensation behavior depends on the magnitude of attraction, as discussed in Fratini and Pieri (2012). Considering the kinetic motion of bosons is strongly correlated with that of fermions, the resulting effective composite-particle Hamiltonian takes the form
$H=$ $-t_{f}\sum_{i}(c^{\dagger}_{i}c_{i+1}+H.c.)+\frac{U_{bb}}{2}\sum_{i}n^{b}_{i}(n^{b}_{i}-1)$ (2)
$-t_{cf}\sum_{i}(c^{\dagger}_{i}b^{\dagger}_{i}c_{i+1}b_{i+1}+H.c.)$
where $t_{cf}$ is the amplitude of composite-particle hopping. Prior to collapse, bosons transition into a superfluid state for a nonzero $t_{cf}$ values. As depicted in Fig. 1(c-d), DMRG calculations for this composite-particle model reveal that the momentum distributions of bosons exhibit analogous even-odd disparities observed in the BFH Hamiltonian(1). Notably, the Fermi surfaces of both fermions and CFs are discernible in both models. (Details of phase diagrams for the two Hamiltonians are given in the Supplementary MaterialSM .)
Origin of CSF from the composite-particle model.— Distinct from DMRG simulations, which maintains a conserved particle count, GMERA operates within the grand-canonical ensemble framework, enabling direct measurement of the local superfluid order parameter $\langle b_{i}\rangle$. In contrast to conventional superfluids, where condensation occurs at $k=0$, non-zero momentum condensation of the CSF leads to a complex local order parameter $\langle b_{i}\rangle$ that rotates continuously along the chain. The twisted angles $\theta_{i,i+1}$ (argument difference between $\langle b_{i}\rangle$ and $\langle b_{i+1}\rangle$ on two neighboring sites) add up to $2\pi$ over a complete lattice period, as illustrated in Fig. 2. This chiral superfluid state spontaneously breaks time-reversal symmetry, resulting in a unique and intriguing emergent flux state in the Bose-Fermi mixture.
We address the origin of CSF from two distinct perspectives, starting from the bare fermion limit ($t_{f}\gg t_{cf}$). First, the kinetic term of composite particles ($t_{cf}$) suggests that a bosonic superfluid in the background induces projected ”silhouette” composite particles. A CSF introduces an additional phase factor that fermions acquire during hopping. Effectively, this generates an ”emergent flux” that further improves the system energy, analogous to the behavior of electrons hopping in a ring with a magnetic fluxPeter et al. (1989). Based on this argument, we develop a mean-field consideration, resulting in an effective boson-free model that qualitatively aligns with our original Hamiltonian $H$ in the bare fermion limit. Second, the fermionic response to the bosonic emergent flux is crucial. As the CSF emerges, it profoundly modifies the fermionic momentum distribution. Numerical calculations reveal that the degeneracy of the fermionic Fermi surface (two Fermi points) is lifted simultaneously when the time-reversal symmetry of the superfluid is broken. This lifting of degeneracy is accompanied by a redistribution of the momentum distribution of composite particles, which becomes more concentrated around small $k$ values, further optimizing the system’s energy. The collaborative effect between the fermionic and bosonic sectors stabilizes the CSF state.
@@
CSF and Fermions parity.— Furthermore, our system exhibits an intriguing sensitivity to the parity of fermions. Specifically, when the number of fermions $N_{f}$ is even, the ground state is in the CSF phase where bosons condensation occurs at $k_{0}=\pm 2\pi/L$, either clockwise or counterclockwise, depending on the sign of $k_{0}$. However, when $N_{f}$ is odd, bosons condense at $k=0$, similar to a normal superfluid. In both cases, the superfluid phase persists to an intense competition region $t_{f}\approx t_{cf}$. Numerical results from the long-range boson correlation $G_{b}(x)=\langle b_{i}^{\dagger}b_{i+x}\rangle$ and momentum distribution $n_{b}(k)$ show that superfluidity emerges as soon as $t_{cf}$ is turned on, such as ($t_{cf}=10^{-3}t_{f}$). In the weakly coupled case, with a boson superfluid as the background, we understand that the free hopping fermions can gain additional energy through the $t_{cf}$ term in Eqn. 2, by projecting out silhouette composite particles. When $N_{f}$ is even, the energy contribution from the $t_{cf}$ term can be further optimized by allowing superfluid order parameter $\langle b\rangle$ to rotate and acquire chirality. In such cases, the hopping amplitude $\langle b_{i}^{\dagger}b_{i+1}\rangle$ becomes a complex number due to the phase difference of $\langle b\rangle$ between neighboring sites. Consequently, the expression $c_{i}^{\dagger}c_{i+1}b_{i}^{\dagger}b_{i+1}$ should be regarded as $c_{i}^{\dagger}c_{i+1}\langle b_{i}^{\dagger}b_{i+1}\rangle$. Numerical results confirm that the measurement of $\langle b_{i}^{\dagger}b_{i+1}\rangle$ on different bond ($i,i+1$) exhibit nearly identical arguments and moduli, as shown in Fig. 2(d). This observation suggests that a CSF generates an additional phase factor $e^{i\theta_{i}}$ in the composite-particle hopping term.
@@
Figure 3: The ground state energy $E(\theta)$ of the mean-field Hamiltonian $H_{MF}$ as a function of the argument $\theta$ in $L=18,36,54$ systems when $t^{\prime}_{cf}=2t_{f}$. (a) Even fermion parity. The fermion filling factor is fixed at $n_{f}=N_{f}/L=4/9$. There is a prominent peak on $\theta=0$ where no magnetic flux is threaded through the system. And $E(\theta)$ reaches its absolute minimum at $\theta_{opt}=\pi/L$. (b) Odd fermion parity. $N_{f}$ are $7,15,23$ for $L=18,36,54$ respectively. No peak shows on $\theta=0$ contrary to the even case in the main panel. And $E(\theta)$ reaches its absolute minimum at $\theta_{opt}=0$
@@
Figure 4: Momentum distribution functions $n(k)$ for $L=36$ system with boson filling factor $n_{b}\approx 1/4$, obtained by GMERA on the composite-particle Hamiltonian(2). ”FS” refers to the symmetric Fermi surface of original fermions. (a) For $N_{f}=16$ (even), bosons condensate at $|k_{0}|=2\pi/L$ (position of the blue peak in $n_{b}(k)$). $k^{c}_{f}=\sum_{k}n_{f}(k)\times k/N_{f}=-0.0122\pi$ is the center of original Fermi sea, whereas $k^{c}_{cf}=\sum_{k}n_{cf}(k)\times k/N_{cf}=0.0075\pi$ is the center of Fermi sea for composite fermion, where $N_{cf}=\sum_{k}n_{cf}(k)$. (b) For $N_{f}=15$ (odd), bosons condensate at $k=0$. The parameters are chosen as $t_{cf}=2t_{f},U_{bb}=7.5t_{f}.$ Fourier transforms of $G_{b}(i,j)=\langle b_{i}^{\dagger}b_{j}\rangle$, $G_{f}(i,j)=\langle c_{i}^{\dagger}c_{j}\rangle$, $G_{cf}(i,j)=\langle c_{i}^{\dagger}b_{i}^{\dagger}c_{j}b_{j}\rangle$ and $G_{hf}(i,j)=\langle c_{i}^{\dagger}b_{i}c_{j}b_{j}^{\dagger}\rangle$ give respective momentum distributions $n_{b}(k),n_{f}(k),n_{cf}(k)$ and $n_{hf}(k)$. The black dashed lines guide the eyes at $k=0$.
@@
$H_{MF}=-t_{f}\sum_{i}(c^{\dagger}_{i}c_{i+1}+H.c.)-t^{\prime}_{cf}\sum_{i}(c^{\dagger}_{i}c_{i+1}e^{i\theta}+H.c.)$
@@
The second term can be intuitively understood through our mean-field ansatz, treating the influence of bosons classically as a pure phase factor $e^{i\theta}$. Therefore, we substitute the hopping term $c_{i}^{\dagger}b_{i}^{\dagger}c_{i+1}b_{i+1}$ with $c_{i}^{\dagger}c_{i+1}e^{-i\theta}$. This model highlights the two competing hopping processes: the former stands for original free hopping, whereas the latter depicts hoppings experiencing a lattice flux. This simple model can be solved straightforwardly. As shown in Fig. 3, the ground state energy $E$ exhibits a periodically varying function superimposed on a parabolic background, illustrating the competition between two distinct hopping processes. Remarkably, a significant difference is also observed between systems with even (left panel) and odd (right panel) fermion parity: when $N_{f}$ is even, the optimal $\theta$ equals to $\pi/L$. In contrast, when $N_{f}$ is odd, the optimal $\theta$ takes the value of 00. This observation from the mean-field model is similar to the results obtained in the BFM. However, it is crucial to note that $\theta_{i}$, the argument of the complex $\langle b_{i}^{\dagger}b_{i+1}\rangle$ shown in Fig. 2(d), averages out to $0.0940\pi$ for $L=18$ or $-0.0491\pi$ for $L=36$, deviating from the optimal angle of ±$\pm\pi/L$ predicted by mean-field theory. This deviation is attributed to the homogeneous condition of our 1D periodic lattice.
@@
Shift of Fermi surface.— Interplay and correlations between different species of particles lead to the shift of the Fermi surface. We focus on the correlation function and momentum distribution of involved particles, influenced by the effective flux. When $N_{f}$ is odd, the conventional bosonic superfluid serves as a flux-free background, the correlation function of projected composite fermions $G_{cf}(i,j)=\langle c_{i}^{\dagger}b_{i}^{\dagger}c_{j}b_{j}\rangle$ shows similar long-range behavior as that of an original fermion. As demonstrated in Fig. 4, two clear Fermi surfaces (points) of composite particles have the same momentum as those of spinless fermions. In contrast, when $N_{f}$ is even, the ground state exhibits a two-fold degeneracy: bosons condensate at either $+2\pi/L$ or $-2\pi/L$, time-reversal symmetry of the system is spontaneously broken. Intriguingly, this degeneracy lift simultaneously affects the Fermi surface of a bare fermion. The selection of the last filled momentum is determined by the direction of the effective flux, resulting in an asymmetric bulk of the Fermi surface. Consequently, the Fermi surface of the projected composite fermion shifts towards $k=0$, optimizing the system’s overall energy. In this uncorrelated limit, this projection can be rationalized as follows: $n_{cf}(k)=1/L\sum_{i,j}G_{cf}(i,j)e^{ik(x_{i}-x_{j})}\approx 1/L\sum_{i,j}G_{f}(i,j)G_{b}(i,j)e^{ik(x_{i}-x_{j})}\approx 1/L\sum_{i,j}G_{f}(i,j)|A|e^{i(k+k_{0})(x_{i}-x_{j})}.$ Here, $G_{b}(i,j)=|A|e^{ik_{0}(x_{i}-x_{j})}$ represents the boson correlation for superfluid condensates at momentum $k_{0}$.
@@
Figure 5: The entanglement entropy $S_{vN}$ as a function of $L$ at different values of $t_{cf}/t_{f}$ and fermion parity in the semilogarithemic plot, obtained by DMRG. Boson and fermion numbers are $N_{f}=2N_{b}=L/2$ in ’a’ (blue solid squares) and ’b’ (red solid circles), $N_{f}=L/2-1,N_{b}=L/2$ in ’c’ (orange solid diamonds) and $N_{f}=N_{b}=L/2$ in ’d’ (green solid triangles). Inset shows the results of central charge.